OpenStudy (theedwardsfamily):

Forty-eight is what percent of 60?

OpenStudy (theedwardsfamily):

@Godlovesme

OpenStudy (anonymous):

what do u think

OpenStudy (anonymous):

80%

OpenStudy (amorfide):

when you deal with percentages let's say we want 20% of 60 since percentages are out of 100 we can write our question like this \[\frac{ 20 \times 60 }{ 100 }=12\] since 12 is 20% of 60 so you want the percentage of 60 that gives 48 so \[\frac{ 60 \times x }{ 100 }=48\] work out x to get the percentage

OpenStudy (anonymous):

48 / 60 x 100 = 80%